OpenStudy (anonymous):

I AM CERTAIN THAT NO ONE WILL ANSWER THIS QUESTION... Suppose that l be a line through the points (0,3)and (5,-2) and tangent to the graph of the function \[F(x)=\frac{ x^2-p }{ 3x+1 }\] where ‘p’ is real number. What is the equation of a line parallel to l and tangent to the graph of F. A) 4x+4y+7 = 0 B) 9x+9y+37 = 0 C) x+y-3 = 0 D) 4x+6y-37 = 0

OpenStudy (anonymous):

I will post the answer exactly after one hour.

mathslover (mathslover):

Hint: Never challenge OpenStudy

OpenStudy (anonymous):

m= -2-3/5-0 = -1 the equation will be y-3= -1(x-0) x+y-3 =0

OpenStudy (theviper):

Right @mathslover :)

hartnn (hartnn):

so its not D

OpenStudy (anonymous):

@hartnn that means???

hartnn (hartnn):

that means answer can be a,b or c

OpenStudy (anonymous):

yep.... but the answer hasn't yet finished... @hartnn

OpenStudy (anonymous):

I will the slope of the line through the two points

hartnn (hartnn):

p=-7 B)9x+9y+37=0

hartnn (hartnn):

solve y=3-x=F(x)=.... u get two roots of x, equate them because the line is tangent and can have only one intersection point. from this i got p=-7. now answer can only be either A or B so solving both of them with F(x), gave only B) option with one intersection point, A) doesn't even intersect.

OpenStudy (anonymous):

PLZ DONT POST THE SOLUTION

OpenStudy (anonymous):

I am here to post the answer.

OpenStudy (anonymous):

ok 5min

OpenStudy (anonymous):

ok

hartnn (hartnn):

whats wrong with my answer? and method ??

OpenStudy (anonymous):

w/c is your ans?

hartnn (hartnn):

B)

hartnn (hartnn):

and p=-7

OpenStudy (anonymous):

correct but the approach not.

hartnn (hartnn):

didn't u get p=-7, the same way as i got ?

OpenStudy (anonymous):

yape

hartnn (hartnn):

then ?

OpenStudy (anonymous):

you use the options to answer the question, isn't it?

OpenStudy (anonymous):

Time is up @sauravshakya

OpenStudy (anonymous):

OK.......

OpenStudy (anonymous):

Here is the answer

OpenStudy (anonymous):

First of all what you have to do is finding l so as to get p \[l:\frac{ y-3 }{ x-0 }=\frac{ -2-3 }{ 5-0 }\] \[l:y=-x+3\]thus \[F(x)=\frac{ x^2-p }{ 3x+1} =-x+3=y\]simplifying the above expression \[4x^{2}-8x-p-3=0\]Now since l is tangent to F(x), the quadratic equation above has only one solution. That means \[b^{2}-4ac=0\] \[(-8)^{2}-4(4)(-p-3)=0\]thus \[(-8)^{2}-4(4)(-p-3)=0\]\[p=-7\]Now we get F(x) to be \[F(x)=\frac{ x^2+7 }{ 3x+1 }\]Therefore lets solve F’(x)=-1 so as to get the coordinate of the unknown line. \[F'(x)=\frac{ 2x(3x+1)-3(x^{2}+7) }{ (3x+1)^2 }=-1\] \[3x^{2}+2x-21=-9x^{2}-6x-1\] \[12x^{2}+8x-20=0\]thus \[x=1 or x=\frac{ -5 }{ 3 }\]and \[F(1)=2 and F(\frac{ -5 }{ 3 })=\frac{ -22 }{ 9 }\]Thus we got this two points (1,2) and (-5/3,-22/9). The two possible lines therefore are \[\frac{ y-2 }{ x-1 }=-1{\rightarrow}y=-x+3\]w/c is the given line l. And \[\frac{ y-(-\frac{ 22 }{9 }) }{ x-(-\frac{ 5 }{ 3 })}=-1{\rightarrow}9y+9x+37=0\]w/c is the answer.

hartnn (hartnn):

why f'(x) = -1 ?

OpenStudy (anonymous):

b/c the slope at that point is -1

OpenStudy (anonymous):

Because the line l and the required line are parallel

OpenStudy (anonymous):

So, both of their slope is -1

hartnn (hartnn):

okk.

hartnn (hartnn):

it wasn't that difficult.

OpenStudy (anonymous):

yes but need smart approach

OpenStudy (anonymous):

YEP....... Actually I did a algebra mistake, as usual

hartnn (hartnn):

lol! checking options(when given) is the smartest approach!! :P

OpenStudy (anonymous):

agreed with @hartnn

OpenStudy (anonymous):

that saves time in exam

OpenStudy (anonymous):

For exam only

hartnn (hartnn):

atleast i got the answer...

OpenStudy (anonymous):

yape

OpenStudy (anonymous):

ok thanks for your help....... closing